# A Review On Safe Reinforcement Learning Using Lyapunov and Barrier Functions

### v2

@@ 4.1 Lyapunov Functions for Reinforcement Learning
Table 1 summarizes the literature reviewed for use of Lyapunov functions in RL. It can be observed that two formal approaches to incorporate Lyapunov functions with RL involve: (i) Lyapunov-based reward-shaping; (ii) CLF-RL-QP formulations; (iii) candidate Lyapunov function as the value/critic function; (iv) Lyapunov/CLF-based loss function. The key challenge lies in approximating or finding a suitable candidate Lyapunov function without having prior or imperfect information of the model/environment. Using model-based methods can circumvent this problem by providing a candidate Lyapunov function based on system dynamics, but its suitability for model-free approaches remains an open area of research. Another difference in methodology for these approaches depend on approximating the Lie derivative of Lyapunov function along the system dynamics, which requires some prior knowledge or is approximated in a data-driven format. Lie derivatives are critical in computing the gradient of loss function with Lyapunov/CLF terms.
4.2 Barrier Functions for Reinforcement Learning
@@
Marvi et al.(2021) [91] Barrier term in reward function fo RL ADP Lane-keeping
Ohnishi et al. (2019) [83] Barrier certified policy updates SARSA Robotarium
Cheng et al. (2019) [26] CBF-based controller compensation TRPO, DDPG Inverted pendulum, car following
Emam et al. (2021) [96] Robust CBF-RL-QP optimization of policy SAC Robotarium
Ma et al. (2021) [97] CBF constraints in policy optimization Actor-critic SUMO
Cai et al. (2021) [98] Decentralized CBF shielding Multi-agent DDPG Particle-env
Zhao et al. (2022) [85] Barrier certificate constraints as a policy optimization problem DDPG OpenAI Gym
Zhang et al. (2022) [84] Barrier terms in critic and policy loss functions Actor-critic Formation Control - mobile robots
Wang et al. (2022) [101] Soft barrier constraints for CBF-RL QP formulation. PPO CommonRoad-RL
Huang et al. (2022) [102] CBF-RL sum of squares programming problem DDPG OpenAI Gym
Song et al. (2022) [103] CBF-RL-QP problem with online learning for CBF SAC ROS Gazebo
@@
Cheng et al. (2023) [104] Disturbance observer with robust CBF-QP policy optimization. SAC Unicycle, quadrotor
@@
Wang et al. (2023) [107] generative-model-based soft barrier functions as constraints in a bilevel policy optimization problem PPO Cartpole, rocket landing
@@
Liu et al. (2023) [88] Discrete exponential CBF constraints in value iteration. Value Iteration Two and three player system
@@
Liu et al. (2023) [92] Discrete barrier constraints and barrier term in reward function using policy iteration Policy Iteration and Actor-critic Inverted pendulum
@@
Yang et al. (2023) [89] Jointly learning barrier certificates and safe policy using NNs. Barrier loss function as constraint on policy PPO Safety Gym, Metadrive
@@
Zhao et al. (2023) [90] Barrier certificate term in policy loss. Formal verification of barrier certificates Actor-critic Power systems
@@
Ranjan et al. (2024) [93] CBF based reward shaping TD3 OpenAI Gym, Unitree Go1 quadruped
@@
Sabouni et al. (2024) [108] Bilevel optimization using MPC-CBF and RL to learn optimal parameters Actor-critic and MPC Multi-agent CAV
Hou et al. (2024) [105] Disturbance observer based CBF-RL QP optimization SAC Gymnasium
@@
Yang et al. (2024) [99] Learning CBF from human demonstration. Using CBF-based shielding in inverse RL. Inverse RL Demolition Derby and Panda Arm Push, real-robot
@@
Zhang et al. (2024) [86] Augment actor loss with smooth log barrier function. SAC Safety Gym, Unitree A1 robot
@@
Dey et al. (2024) [87] Controllable soft barrier penalty term in policy optimization. PPO Safety Gym, Safe MuJoCo
@@
Dinh et al. (2024) [100] CBF-based filter for policy. CBF obtained using local search for constraint violation. PPO, DDPG OpenAI Gym
@@
Kalaria et al. (2024) [106] residual model and Disturbance observer based CBF-RL QP optimization PPO Safety Gym, F1/10 racing car
@@
Liu et al. (2025) [94] Robust neural CBF based reward shaping. MADDPG Custom - robot navigation
Xie et al. (2025) [95] CBF-based reward shaping. AC Gymnasium, HoloOcean
@@
There has been significant recent research on use of barrier functions with RL to define safe sets for the control policy. Fig. 5 shows a broad classification for the use of barrier function-based approaches in safe RL. Similar to reward-shaping using CLF/Lyapunov terms in Section 4.1, reward-shaping using CBF terms is observed in several works to ensure state constraints are satisfied. Marvi et al. [91] propose adding a CBF term to the cost-to-go function for a safe optimal control problem with nonlinear dynamics. A coefficient $\gamma$ (damping term) is introduced along with the CBF term to control the effect of barrier term, this can be interpreted as a trade-off between safety and optimality. The authors also prove the additional barrier term maintains safety and optimality (under the assumption of finite value barrier function and suitable value of $\gamma$). The MPC formulation for a discrete time control system is augmented as follows:
@@
where, $c_{0}(\cdot)$, $c(\cdot)$ and $c_{f}(\cdot)$ are the initial, stage and terminal cost, respectively. $B_{x}(\cdot)$, $B_{u}(\cdot)$ and $B_{f}(\cdot)$ are barrier functions for constraints on states, control input and terminal conditions. An actor-critic algorithm is used to prove efficacy of the proposed approach for a lane-keeping problem. Authors in [92] propose a similar reward-shaping where the reward function is augmented with an exponential CBF term. A policy iteration (PI) and actor-critic RL framework is considered for an inverted pendulum environment. Reward-shaping is further explored in [93, 94, 95] considering exponential, quadratic and robust neural CBFs.
In [83], authors propose a barrier-certified adaptive RL for a quadrotor control and brushbot navigation application using Gaussian process (GP) SARSA algorithm. The proposed approach follows a few key steps: (i) Adaptive model learning (ii) Structured model learning using sparse optimization satisfying monotone approximation property (Section III-B of cited work) in a defined reproducible kernel Hilbert space (RKHS), to capture the structure of agent dynamics (useful for computing Lie derivative of control barrier functions); (iii) Adaptive action-value function approximation with barrier-certified policy updates (Section IV-A). The authors prove global optimality of solutions to update the control policy. Cheng et al. [26] propose a policy combining model-free RL and model-based CBF. The algorithm is tested on inverted pendulum and car-following with vehicle-to-vehicle communication using TRPO and DDPG based RL. The authors propose two approaches combining CBF and RL controller: (i) The first uses a concept similar to CBF shielding (only compensates for RL controller in case of constraint violation and formulated by solving a QP at each iteration) given as,
@@
In this case, dependence on all prior CBF controller enhances learning efficiency. The second approach guides exploration as well by taking prior computed CBFs into consideration. Probabilistic guarantees are provided for TRPO and DDPG algorithms implemented. In [97], authors propose a generalized CBF for a model-based constrained RL policy optimization algorithm. The problem is formulated for an actor-critic algorithm with the actor policy update as an optimization problem with CBF constraints. Lagrange multipliers are used to reformulate the problem as an unconstrained optimization problem along with a distance constraint (collision avoidance simulations). Furthermore, a parameter $\xi$ (constraint violation metric) is computed to compare with a predetermined threshold for updating the parameters of CBF to adjust constraint satisfaction. A robust CBF is proposed in [96] for a nonlinear system with disturbance. The disturbance is approximated using a GP (this approximated model is used to compute the Lie derivatives in (12) and the robust CBF is combined with RL actions to form a relaxed QP problem. The slack variable is introduced to avoid deadlock when no safe action is available. The proposed approach is tested on unicycle and car following environments using a SAC-Robust CBF formulation. Similar optimization formulations are observed in [102, 103, 104, 99, 100, 106]. Soft-barrier constraints are proposed in [101], where a slack variable $\gamma$ is introduced to relax the barrier constraints and ensure a feasible solution for the CBF-RL-QP formulation. A multi-agent RL with decentralized CBF shields is proposed by authors in [98] for a collision-avoidance problem. The two types decentralized CBF shields are formulated namely, cooperative and non-cooperative, the safe policy is computed as a constrained optimization problem for the two types of CBF shields and action by the RL agent, similar to (20). A multi-agent DDPG-CBF formulation is tested on a two patrolman task. In [104], authors use disturbance observers and CBFs with RL to provide safety guarantees. A policy is sampled from the RL agent followed by disturbance estimation from the disturbance-based observer, similar to prior approaches discussed Lie derivative is approximated using disturbance-based observers. A QP (20) is solved using the CBF as constraints and the action sampled by the RL agent to solve for the safe action. SAC algorithm is used for simulations on a 2D quadrotor and a unicycle model. Hou et al. [105] and Kalaria et al. [106] follow a similar disturbance-observer-based CBF filter enforce hard constraints.
@@
Similar to CLF/Lyapunov violations considered in policy loss function, CBF terms are augmented to the loss functions in barrier-based safe RL approaches. Zhao et al. [85] propose a learner-verifier framework where barrier terms are added to the actor loss function and further actions taken by the agent are formally verified by solving an optimization for each barrier certificate. Authors in [84, 88, 89, 87, 86] follow the Lagrangian method to incorporate the barrier constraints in the actor objective function.
@@
Authors in [107] use soft barrier constraints through generative-model-based RL in an unknown stochastic environment. Their approach forces hard safety chance constraints (probabilistic safety constraints) using bilevel optimization formulation. The lower problem encodes hard safety constraints with a generative-model-based soft barrier function and the upper problem maximizes the total expected return of the policy. One limitation of this approach is the high computational complexity of the generative model, as stated by the authors. Sabouni et al. [108], follow a bilevel optimization architecture as well. However, RL is used to learn the optimal parameters of a MPC formulation.

### v3

@@ 4 Lyapunov Functions for Reinforcement Learning
4.1 Classification and Discussion
Table 2 provides a tabular comparison of each methodology and its shortcomings. The body of work surveyed in Section 4 demonstrates that Lyapunov functions provide one of the most direct bridges between classical stability theory and modern reinforcement learning. Broadly, existing approaches use Lyapunov structure in three ways: (i) as a certificate for stability/safety (e.g., ROA/level-set reasoning), (ii) as a constraint that shapes or restricts policy updates (e.g., CMDP feasibility, projection, shielding), and (iii) as a learning signal (e.g., penalties, reward shaping, or empirical risk minimization). These perspectives yield complementary strengths, but they also expose recurring limitations that motivate several key research directions.
Stability certificates: constructing and maintaining valid Lyapunov functions
A central theme is to interpret a Lyapunov candidate as a stability certificate that induces attractive and invariant level sets. Early work uses this certificate to ensure safe switching among stabilizing control laws while still permitting RL to optimize high-level decisions [perkins2002lyapunov]. A more common modern pathway is to identify the value function (or $Q$-function) with a Lyapunov candidate, leveraging approximate dynamic programming and actor–critic updates with Lyapunov analysis [vamvoudakis2015asymptotically, kamalapurkar2016model, kumar2017diagonal, yao2025lyapunov]. While this creates a principled connection to optimal control (e.g., HJB structure), the resulting guarantees often rely on restrictive assumptions about approximation error, excitation conditions, or the dynamics class (frequently deterministic or structured nonlinear systems). In practice, the learned critic may satisfy Lyapunov-like behavior only on the data distribution encountered during training, raising concerns about certificate validity under distribution shift and robustness to function approximation error.
Safe-set expansion with learned models and uncertainty: conservatism vs scalability
A second line of work explicitly learns models and uses Lyapunov level sets to define a safe region of attraction. The GP-based formulation in [berkenkamp2017safe] exemplifies this idea: uncertainty-aware upper bounds are used to enforce a Lyapunov decrease condition over a discretized state set, and the safe region is expanded as data accrue. Related methods incorporate empirical Lyapunov risk into learning objectives [xia2024estimating]. These approaches provide a tangible notion of “where the agent is safe”, but they also surface two practical bottlenecks: (i) the frequent need for an initial safe policy or safe set (otherwise exploration cannot be certified), and (ii) conservatism induced by Lipschitz bounds, discretization, or uncertainty envelopes. Moreover, many uncertainty quantification mechanisms (e.g., GP kernels) can be hard to scale and tune in high-dimensional state spaces, which remains a barrier to deployment in complex robotic domains.
CMDP and constrained optimization: feasibility, projections, and bias
Compared to control-affine stability analyses, Lyapunov-based CMDP formulations provide a unifying RL interface for safety constraints. The Lyapunov-constrained CMDP viewpoint in [chow2018lyapunov] introduces a Bellman-style feasibility condition and yields safe dynamic programming and RL algorithms; extensions include continuous-action policy optimization via parameter projection ($\theta$-projection) and action projection ($a$-projection) [chow2019lyapunov]. Closely related ideas use Lyapunov drift constraints in a QP-like projection of the RL action [cao2023toward], and shielding mechanisms construct safe policies and exploration strategies using Lyapunov-induced safe sets [huh2020safe]. While these methods can enforce hard safety at execution time, they also introduce new failure modes: the feasible set may be empty or numerically fragile, projections may bias gradients and slow convergence, and aggressive constraints can suppress exploration, often yielding safe but overly conservative policies.
Table 2: Classification of Lyapunov-function-based safe RL.
Class Methodology Works Pros Cons
Switching CLF controllers RL switches among stabilizing modes [perkins2002lyapunov] Safety inherits from each mode; faster learning. Needs library of stabilizers; limited expressivity.
Value-as-Lyapunov (ADP/HJB) Use $V$ or $Q$ as Lyapunov candidate [vamvoudakis2015asymptotically, kamalapurkar2016model, kumar2017diagonal, yao2025lyapunov] Tight link to optimal control; stability proofs under assumptions. Strong assumptions; approximation error sensitivity; constraints need extra handling.
Reachability CLVF CLVF/safe set via HJR reach-avoid [lopez2024decomposing] Strong constraint semantics; interpretable sets. Poor scaling with dimension; conservative sets.
Model+ROA expansion (uncertainty) Learn model (e.g., GP), enforce Lyapunov decrease over ROA [berkenkamp2017safe, xia2024estimating] Explicit safe region; can expand with data. Often needs safe initializer; Lipschitz/discretization conservatism; GP scaling.
CMDP Lyapunov feasibility Lyapunov constructs feasible set in CMDP [chow2018lyapunov, chow2019lyapunov] RL-native constrained framework; DP/PG variants. Feasibility brittleness; conservative; projection bias.
Projection / shielding layers Project action/params to satisfy Lyapunov constraints [chow2019lyapunov, cao2023toward, huh2020safe] Harder safety at execution; modular. QP/projection cost; feasibility; reduced exploration.
Reward shaping (Lyapunov) Add Lyapunov temporal-difference shaping [ng1999policy, dong2020principled, ugurlu2025lyapunov] Simple; can accelerate learning. Usually soft safety; weight tuning; can distort optimality.
Self-learned Lyapunov critics Learn $V_{\theta}$ via empirical Lyapunov risk [liu2020almost, chang2021stabilizing, mccutcheon2025neural, hao2024lyapunov] Model-light; adaptive certificates. Drift estimation hard; noisy finite-differences; OOD generalization risk.
Physics-informed / Zubov critic Critic structured as maximal Lyapunov/Zubov [wang2024actor, zubov1964methods] Strong inductive bias; robust ROA reasoning. Coupled/stiff training; approximation reliance.
Two-stage CLF + violation loss Learn CLF then penalize violations in RL [10155918] Simple pipeline; reduces unsafe rollouts after CLF. Soft safety unless filtered; compounding errors.
Multi-agent Lyapunov drift Drift penalties in MARL objectives [krishna2023finite, zhuang2019sdn] Extends to distributed settings. Non-stationarity; coupled constraints; weaker guarantees.
Learning signals: penalties, shaping, and self-learned Lyapunov critics
Several works relax hard constraints and instead use Lyapunov structure as a learning signal. Reward shaping methods add Lyapunov-inspired temporal difference terms to accelerate learning [dong2020principled, ugurlu2025lyapunov, ng1999policy]. These techniques are attractive due to their simplicity and compatibility with standard deep RL pipelines, but safety becomes “soft”: violations may still occur, and performance is sensitive to shaping weights and discounting assumptions. A more ambitious direction is to learn the Lyapunov function itself from data by minimizing empirical Lyapunov risk under relaxed (“almost”) Lyapunov conditions [liu2020almost, chang2021stabilizing, mccutcheon2025neural, hao2024lyapunov]. In this family, the key technical difficulty is reliably estimating Lyapunov drift/Lie derivatives in a model-free setting, where finite-difference approximations can be noisy and biased. As a result, stability assurances may be empirical or probabilistic, and learned certificates can generalize poorly beyond the training distribution.
Physics-informed and maximal Lyapunov structures
Recent work explores injecting stronger inductive bias into the critic by adopting physics-informed structures such as Zubov/maximal Lyapunov functions [wang2024actor, zubov1964methods]. These approaches aim to approximate “largest” regions of attraction and iteratively update actor and critic to satisfy Lyapunov conditions. Although promising for robustness and interpretability, such methods can lead to stiff coupled optimization (actor improvement must not destroy certificate validity), and success can depend strongly on how Lyapunov constraints are enforced during learning.
Extensions: inaccurate models, two-stage pipelines, and multi-agent settings
Beyond standard single-agent settings, CLF dissipation constraints with imperfect models provide a pragmatic compromise: stability is encouraged/enforced through minimum dissipation constraints while using model-free policy optimization [westenbroek2020learning, westenbroek2022lyapunov]. Two-stage pipelines that first learn a CLF (and possibly dynamics) and then penalize Lyapunov violations during RL training have also been reported [10155918]. Finally, Lyapunov drift has been incorporated into multi-agent actor–critic objectives to stabilize distributed learning and coordination [krishna2023finite, zhuang2019sdn], though non-stationarity and credit assignment complicate both certificate learning and theoretical analysis.
4.2 Key Research Issues and Open Problems
Reducing dependence on an initial safe policy or safe set
Many provably safe schemes assume a known safe baseline (or a seed ROA) to bootstrap learning [berkenkamp2017safe, chow2019lyapunov]. Developing safe exploration without a certified initializer remains a critical gap, particularly for real robots where resets and failures are costly.
Certificate validity under function approximation and distribution shift
Deep function approximators can produce Lyapunov candidates that appear stable on-policy but fail off-policy. Robust training objectives, explicit generalization tests for certificates, and distributionally robust formulations for Lyapunov decrease constraints are needed, especially in high-dimensional continuous control [chang2021stabilizing, westenbroek2020learning].
Reliable estimation of Lyapunov drift/Lie derivatives in model-free settings
Finite-difference approximations used in self-learned Lyapunov critics can be sensitive to noise and step size [chang2021stabilizing]. Open problems include variance reduction, uncertainty-aware drift estimation, and hybrid approaches that exploit partial models or learned latent dynamics while maintaining valid bounds [berkenkamp2017safe].
Conservatism, feasibility, and the safety performance trade-off
Hard constraints via projection/shielding can become overly conservative or even infeasible [chow2019lyapunov, cao2023toward, huh2020safe]. A major practical challenge is designing adaptive constraint tightening/relaxation strategies and diagnosing when conservatism stems from certificate choice, discretization/Lipschitz bounds, or approximation error.
Scaling certificates and uncertainty quantification
Reachability-based constructions yield strong semantics but scale poorly with dimension [lopez2024decomposing]. GP-based uncertainty methods can be difficult to scale and tune [berkenkamp2017safe]. Promising directions include compositional certificates, structured Lyapunov parameterizations, and scalable uncertainty surrogates that preserve conservative bounds.
Unifying CLF/QP-style filters with deep RL training dynamics
Action filters and QPs can guarantee safety during execution, but their interaction with policy gradients can introduce bias and learning instability [chow2019lyapunov, cao2023toward]. Better theory and practice are needed for end-to-end training with differentiable safety layers and for quantifying how projection affects optimality and convergence.
Multi-agent and partially observed environments
Lyapunov drift augmentation in MARL is promising [krishna2023finite], but certificates become harder to define when agents induce non-stationary dynamics or when safety is coupled across agents. Incorporating memory and uncertainty models (e.g., Transformer-based risk prediction) in CMDPs [jeddi2021lyapunov] suggests a path forward, but scalable theory for partial observability and joint safety constraints is still underdeveloped.
Table 1 summarizes the literature reviewed for use of Lyapunov functions in RL. It can be observed that four formal approaches to incorporate Lyapunov functions with RL involve: (i) Lyapunov-based reward-shaping; (ii) CLF-RL-QP formulations; (iii) candidate Lyapunov function as the value/critic function; (iv) Lyapunov/CLF-based loss function.
5 Barrier Functions for Reinforcement Learning
@@
Song et al. (2022) [song2022safe] CBF-RL-QP problem with online learning for CBF SAC ROS Gazebo
@@
Cheng et al. (2023) [cheng2023safe] Disturbance observer with robust CBF-QP policy optimization. SAC Unicycle, quadrotor
@@
Wang et al. (2023) [wang2023enforcing] generative-model-based soft barrier functions as constraints in a bilevel policy optimization problem PPO Cartpole, rocket landing
@@
Liu et al. (2023) [liu2023safe1] Discrete exponential CBF constraints in value iteration. Value Iteration Two and three player system
@@
Liu et al. (2023) [liu2023safe2] Discrete barrier constraints and barrier term in reward function using policy iteration Policy Iteration and Actor-critic Inverted pendulum
@@
Yang et al. (2023) [yang2023model] Jointly learning barrier certificates and safe policy using NNs. Barrier loss function as constraint on policy PPO Safety Gym, Metadrive
@@
Zhao et al. (2023) [zhao2023barrier] Barrier certificate term in policy loss. Formal verification of barrier certificates Actor-critic Power systems
@@
Ranjan et al. (2024) [ranjan2024barrier] CBF based reward shaping TD3 OpenAI Gym, Unitree Go1 quadruped
@@
Sabouni et al. (2024) [sabouni2024reinforcement] Bilevel optimization using MPC-CBF and RL to learn optimal parameters Actor-critic and MPC Multi-agent CAV
Hou et al. (2024) [hou2024safe] Disturbance observer based CBF-RL QP optimization SAC Gymnasium
@@
Yang et al. (2024) [yang2024enhancing] Learning CBF from human demonstration. Using CBF-based shielding in inverse RL. Inverse RL Demolition Derby and Panda Arm Push, real-robot
@@
Zhang et al. (2024) [zhang2024constrained] Augment actor loss with smooth log barrier function. SAC Safety Gym, Unitree A1 robot
@@
Dey et al. (2024) [dey2024p2bpo] Controllable soft barrier penalty term in policy optimization. PPO Safety Gym, Safe MuJoCo
@@
Dinh et al. (2024) [dinh2024towards] CBF-based filter for policy. CBF obtained using local search for constraint violation. PPO, DDPG OpenAI Gym
@@
Kalaria et al. (2024) [kalaria2024disturbance] residual model and Disturbance observer based CBF-RL QP optimization PPO Safety Gym, F1/10 racing car
@@
Liu et al. (2025) [liu2025safe] Robust neural CBF based reward shaping. MADDPG Custom - robot navigation
Xie et al. (2025) [xie2025certificated] CBF-based reward shaping. AC Gymnasium, HoloOcean
@@
There has been significant recent research on use of barrier functions with RL to define safe sets for the control policy. Fig. 5 shows a broad classification for the use of barrier function-based approaches in safe RL. Similar to reward-shaping using CLF/Lyapunov terms in Section 4, reward-shaping using CBF terms is observed in several works to ensure state constraints are satisfied. Marvi et al. [marvi2021safe] propose adding a CBF term to the cost-to-go function for a safe optimal control problem with nonlinear dynamics. A coefficient $\gamma$ (damping term) is introduced along with the CBF term to control the effect of barrier term, this can be interpreted as a trade-off between safety and optimality. The authors also prove the additional barrier term maintains safety and optimality (under the assumption of finite value barrier function and suitable value of $\gamma$). The MPC formulation for a discrete time control system is augmented as follows:
@@
where, $c_{0}(\cdot)$, $c(\cdot)$ and $c_{f}(\cdot)$ are the initial, stage and terminal cost, respectively. $B_{x}(\cdot)$, $B_{u}(\cdot)$ and $B_{f}(\cdot)$ are barrier functions for constraints on states, control input and terminal conditions. An actor-critic algorithm is used to prove efficacy of the proposed approach for a lane-keeping problem. Authors in [liu2023safe2] propose a similar reward-shaping where the reward function is augmented with an exponential CBF term. A policy iteration (PI) and actor-critic RL framework is considered for an inverted pendulum environment. Reward-shaping is further explored in [ranjan2024barrier, liu2025safe, xie2025certificated] considering exponential, quadratic and robust neural CBFs.
In [ohnishi2019barrier], authors propose a barrier-certified adaptive RL for a quadrotor control and brushbot navigation application using Gaussian process (GP) SARSA algorithm. The proposed approach follows a few key steps: (i) Adaptive model learning (ii) Structured model learning using sparse optimization satisfying monotone approximation property (Section III-B of cited work) in a defined reproducible kernel Hilbert space (RKHS), to capture the structure of agent dynamics (useful for computing Lie derivative of control barrier functions); (iii) Adaptive action-value function approximation with barrier-certified policy updates (Section IV-A). The authors prove global optimality of solutions to update the control policy. Cheng et al. [cheng2019end] propose a policy combining model-free RL and model-based CBF. The algorithm is tested on inverted pendulum and car-following with vehicle-to-vehicle communication using TRPO and DDPG based RL. The authors propose two approaches combining CBF and RL controller: (i) The first uses a concept similar to CBF shielding (only compensates for RL controller in case of constraint violation and formulated by solving a QP at each iteration) given as,
@@
In this case, dependence on all prior CBF controller enhances learning efficiency. The second approach guides exploration as well by taking prior computed CBFs into consideration. Probabilistic guarantees are provided for TRPO and DDPG algorithms implemented. In [ma2021model], authors propose a generalized CBF for a model-based constrained RL policy optimization algorithm. The problem is formulated for an actor-critic algorithm with the actor policy update as an optimization problem with CBF constraints. Lagrange multipliers are used to reformulate the problem as an unconstrained optimization problem along with a distance constraint (collision avoidance simulations). Furthermore, a parameter $\xi$ (constraint violation metric) is computed to compare with a predetermined threshold for updating the parameters of CBF to adjust constraint satisfaction. A robust CBF is proposed in [emam2021safe] for a nonlinear system with disturbance. The disturbance is approximated using a GP (this approximated model is used to compute the Lie derivatives in (12) and the robust CBF is combined with RL actions to form a relaxed QP problem. The slack variable is introduced to avoid deadlock when no safe action is available. The proposed approach is tested on unicycle and car following environments using a SAC-Robust CBF formulation. Similar optimization formulations are observed in [huang2022barrier, song2022safe, cheng2023safe, yang2024enhancing, dinh2024towards, kalaria2024disturbance]. Soft-barrier constraints are proposed in [wang2022ensuring], where a slack variable $\gamma$ is introduced to relax the barrier constraints and ensure a feasible solution for the CBF-RL-QP formulation. A multi-agent RL with decentralized CBF shields is proposed by authors in [cai2021safe] for a collision-avoidance problem. The two types decentralized CBF shields are formulated namely, cooperative and non-cooperative, the safe policy is computed as a constrained optimization problem for the two types of CBF shields and action by the RL agent, similar to (20). A multi-agent DDPG-CBF formulation is tested on a two patrolman task. In [cheng2023safe], authors use disturbance observers and CBFs with RL to provide safety guarantees. A policy is sampled from the RL agent followed by disturbance estimation from the disturbance-based observer, similar to prior approaches discussed Lie derivative is approximated using disturbance-based observers. A QP (20) is solved using the CBF as constraints and the action sampled by the RL agent to solve for the safe action. SAC algorithm is used for simulations on a 2D quadrotor and a unicycle model. Hou et al. [hou2024safe] and Kalaria et al. [kalaria2024disturbance] follow a similar disturbance-observer-based CBF filter enforce hard constraints.
Table 4: Classification of barrier-function-based safe RL.
Class Methodology Works Pros Cons
CBF reward shaping Barrier terms in reward/cost [marvi2021safe, liu2023safe2, ranjan2024barrier, liu2025safe, xie2025certificated] Simple integration; helps exploration. Soft safety; sensitive weights; may be over-conservative.
Barrier-certified learning (GP/RKHS) Learn structure to compute Lie derivatives; certify updates [ohnishi2019barrier] Uncertainty-aware certificates; strong theory (assumption-heavy). GP/RKHS scaling; conservatism; derivative estimation burden.
CBF-QP shielding (filter) $u^{safe}=\arg\min\|u-u^{RL}\|$ s.t. CBF [cheng2019end, emam2021safe, huang2022barrier, song2022safe, cheng2023safe, yang2024enhancing, dinh2024towards, hou2024safe, kalaria2024disturbance] Hard constraint enforcement when feasible; modular. Needs model/derivatives; QP cost; feasibility/deadlock; exploration restriction.
Soft CBF / slack QP Slack variables to guarantee feasibility [wang2022ensuring, emam2021safe] Avoids infeasibility/deadlock. Safety becomes graded; slack tuning; may hide modeling error.
Constrained policy optimization Actor update constrained by CBF; Lagrange/primal-dual [ma2021model, zhang2022barrier, liu2023safe1, yang2023model, dey2024p2bpo, zhang2024constrained] End-to-end safety-aware learning; can reduce per-step QP. Often expectation/approx. safety; multiplier tuning; non-convexity.
Learner–verifier frameworks Barrier loss + formal verification step [zhao2022safe] Stronger assurance via verification. Verification overhead; scalability depends on verifier.
Robust/disturbance-aware CBFs GP disturbance models/observers for robust CBF-QP [emam2021safe, cheng2023safe, hou2024safe, kalaria2024disturbance] Better under perturbations/unmodeled effects. Estimator error weakens guarantees; added tuning/latency.
Multi-agent decentralized shields Local CBF shields (coop/non-coop) [cai2021safe] Distributed collision avoidance; modular. Coupled feasibility; non-stationarity; network assumptions.
Chance constraints / bilevel Probabilistic safety via bilevel soft barriers [wang2023enforcing, sabouni2024reinforcement] Targets stochastic safety directly. High compute; calibration sensitivity; bilevel complexity.
@@
Similar to CLF/Lyapunov violations considered in policy loss function, CBF terms are augmented to the loss functions in barrier-based safe RL approaches. Zhao et al. [zhao2022safe] propose a learner-verifier framework where barrier terms are added to the actor loss function and further actions taken by the agent are formally verified by solving an optimization for each barrier certificate. Authors in [zhang2022barrier, liu2023safe1, yang2023model, dey2024p2bpo, zhang2024constrained] follow the Lagrangian method to incorporate the barrier constraints in the actor objective function.
@@
Authors in [wang2023enforcing] use soft barrier constraints through generative-model-based RL in an unknown stochastic environment. Their approach forces hard safety chance constraints (probabilistic safety constraints) using bilevel optimization formulation. The lower problem encodes hard safety constraints with a generative-model-based soft barrier function and the upper problem maximizes the total expected return of the policy. One limitation of this approach is the high computational complexity of the generative model, as stated by the authors. Sabouni et al. [sabouni2024reinforcement], follow a bilevel optimization architecture as well. However, RL is used to learn the optimal parameters of a MPC formulation.